OpenStudy (melissa07):

If sin Θ = 2 over 3 and tan Θ < 0, what is the value of cos Θ?

OpenStudy (melissa07):

@whpalmer4 @hahahahazxcvv @TheOcean

OpenStudy (whpalmer4):

A good first step is to draw a diagram showing what you know.

OpenStudy (whpalmer4):

we'll fix up the sign the of the result once we find its magnitude.

OpenStudy (whpalmer4):

if \(\sin \theta = 2/3\) then we know the opposite side and hypotenuse are in that ratio I've drawn. You can use Pythagoras to find the value of \(x\). From that, you can get \(\cos \theta\) but remember that \(x\) will have to be treated as negative